You are given the circle "x^{2}+y^{2}+ax+by+c=0", where a, b and c are some constants. Given that this circle passes through the points (4,2); (0,3) and (3,-2).


Which of the following are true?


  1. The system of equations connecting a, b and c is "\begin{array}{cccc} 4a& +2b& +c&= -20\\ & +3b & +c& =-9 \\ 3a& -2b&+c&=-13 \end{array}"
  2. The determinant of the coefficient matrix for the system connecting a, b and c is 17
  3. With Cramer's rule, the value of a is "a=\frac{\left| \begin{array}{cccc}-20 & 2& 1\\ -9 & 3& 1\\ -13 & -2& 1 \end{array} \right |}{17}."
  4. With Cramer's rule, the value of b is "b=\frac{\left| \begin{array}{cccc}-20 & 2& 1\\ -9 & 3& 1\\ -13 & -2& 1 \end{array} \right |}{17}."
1
Expert's answer
2021-09-16T00:47:00-0400

firstly putting the values of points one by one in the given equation points are  (4,2), (0,3) and (3,-2).

"x^{2}+y^{2}+ax +by +c=0"

putting value (4,2) in equation we get

"16+4+4a+2b+c=0\\4a+2b+c=-20"

putting value (0,3) in equation we get

"0+9+0+3b+c=0\\3b+c=-9"

putting value (3,-2) in equation we get

"9+4+3a-2b+c=0"

"3a-2b+c=-13"

this shows that first option is correct

2 .after finding the determinant of matrix which is 17 second option is also correct

3 . according to cramer's rule the value of a is correct so third option is also correct

4 . fourth option is violating cramer's rule because it has its cofficient of b in the determinant